OpenStudy (anonymous):

Point G is located at (3, -1) and point H is located at (-2, 3). Find the point that is 2/3 the distance from point G to point H. (5 points) Select one: a. (0.33, -1.67) b. (-0.33, 1.67) c. (-0.5, -1) d. (6.33, -2.67)

OpenStudy (anonymous):

@purplepiggy171

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the section formula :- point dividing the segment joining points \((x_1, y_1)\) and \( (x_2, y_2)\) internally in ratio \(m : n\) = \(\large \left( \frac{mx_2+nx_1}{m+n}, \frac{my_2+ny_1}{m+n} \right)\)

OpenStudy (anonymous):

so what would it Be

OpenStudy (pratyush5):